OpenStudy (anonymous):

simplify cos(4x)cos(3x) - sin(4x)sin(3x)

OpenStudy (midhun.madhu1987):

cos A.cos B - sin A.sin B = cos(A+B) Try to do with this formula

OpenStudy (anonymous):

im not sure how lol

OpenStudy (anonymous):

can someone do a walk through ?

OpenStudy (midhun.madhu1987):

cos(4x)cos(3x) - sin(4x)sin(3x) = cos(4x-3x) = cos x

OpenStudy (midhun.madhu1987):

cos A.cos B - sin A.sin B = cos(A+B) A = 4x and B = 3x

OpenStudy (anonymous):

ohhhhh makes sense haha thanks!

OpenStudy (midhun.madhu1987):

:)